# The Trojan Knowledge: Bypassing Commercial LLM Guardrails via Harmless Prompt Weaving and Adaptive Tree Search

### v2

@@ B.1 Performance of ALL Methods on Different LLM Judges
In this section, we present a comprehensive evaluation of all jailbreak methods across multiple judge models to assess the robustness and consistency of our empirical findings (Tab. 4). Beyond Gemini-2.5-Flash (used in the main results), we evaluate two additional powerful commercial judges: Gemini-2.5-Pro and GPT-5-mini. Among these, Gemini-2.5-Pro employs the most stringent evaluation criteria, consistently assigning lower success rates across all methods. Despite differences in absolute values, all three judges exhibit remarkably consistent ranking patterns: CKA-Agent achieves the highest Full Success rates, followed by Multi-Agent Jailbreak, while prompt-optimization methods such as PAIR,

### v3

@@ B.1 Performance of ALL Methods on Different LLM Judges
In this section, we present a comprehensive evaluation of all jailbreak methods across multiple judge models to assess the robustness and consistency of our empirical findings (Tab. LABEL:tab:compare_diff_judge_model). Beyond Gemini-2.5-Flash (used in the main results), we evaluate two additional powerful commercial judges: Gemini-2.5-Pro and GPT-5-mini. Among these, Gemini-2.5-Pro employs the most stringent evaluation criteria, consistently assigning lower success rates across all methods. Despite differences in absolute values, all three judges exhibit remarkably consistent ranking patterns: CKA-Agent achieves the highest Full Success rates, followed by Multi-Agent Jailbreak, while prompt-optimization methods such as PAIR,
@@ Appendix D Details on Jailbreaking Methods
TAP is an automated black-box jailbreak framework that casts prompt generation as a structured tree search with explicit branching and pruning to improve both effectiveness and query efficiency. Building on PAIR, TAP treats each attack as a node in a search tree and uses an attacker LLM to generate multiple refined prompt variants at each iteration, while an evaluator scores candidates for topical relevance and jailbreak likelihood and prunes off-topic or low-quality prompts to control redundancy and cost. In our experiments, TAP is configured with 2 root nodes, a branching factor of 2, a maximum width of 3 nodes per level, and a maximum depth of 3, with early stopping when the evaluator score reaches 8.